OpenStudy (anonymous):

Prove that if a>b>0, then (a−b)/a < ln a − ln b < (a−b)/b.

OpenStudy (anonymous):

if\[ a>b>0\]then\[\frac{ a-b }{ a }>\ln a -\ln b>\frac{ a-b }{ b }\]

OpenStudy (anonymous):

\[a>b>0\implies{a\over b}>1\implies\ln{a\over b}>0\qquad\text{...(1)}\\ a>b>0\implies{b\over a}<1\\ a>b>0\implies a-b>0\implies {a-b\over b}={a\over b}-1>2\qquad\text{...(2)}\\ \rm{\&}\qquad{a-b\over a}=1-{b\over a}<0\qquad\text{...(3)} \] from 1, 2, and 3. QED

OpenStudy (anonymous):

corrections with (2) and (3) it should be the other way... \[{a-b\over b}<0\\{a-b\over a}>2\]

OpenStudy (anonymous):

to be more specific, inequality(1) can be expressed as: \[{a\over b}>\ln{a\over b}>0\]

OpenStudy (anonymous):

but shouldn't\[\frac{ a-b }{ b }=\frac{ a }{ b }-1>0\] if \[\frac{ a }{ b }>1\] and in the same way shouldn't \[\frac{ a-b }{ a }=1-\frac{ b }{ a }>0\] if \[ \frac{ b }{ a }<1\]? @electrokid

OpenStudy (anonymous):

@Bomull lol yeah.. But you got the point to preceed.. :)

OpenStudy (anonymous):

I'm still stuck with this one, if anyone can give me any pointers. How do I know that\[\frac{ a-b }{ a }>\ln a-\ln b\] and\[\ln a-\ln b>\frac{ a-b }{ b }\] ?

OpenStudy (anonymous):

\[\huge alog (\frac{ a }{ b })+b>a\]

OpenStudy (anonymous):

\[\huge b*log \frac{ a }{ b }+b<a\]

OpenStudy (anonymous):

using the 1st and 2nd equations we cann prove it!!

OpenStudy (anonymous):

I don't get it..

OpenStudy (anonymous):

oh ok so if \[a*\ln \frac{ a }{ b }<a\] then\[\ln \frac{ a }{ b }+b<\frac{ a }{ a }\] \[\frac{ a-b }{ a }>\ln \frac{ a }{ b }\]

OpenStudy (anonymous):

but where do I pull off \[a*\ln \frac{ a }{ b }+b<a\]...

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@electrokid @deena123

OpenStudy (anonymous):

\[\ln \frac{ a }{ b }+1-\frac{ b }{ a }>0\] \[\ln \frac{ a }{ b }+\frac{ b }{ a }>-1\] \[\ln \frac{ a }{ b }>\frac{ b }{ a }-1\] \[\ln \frac{ a }{ b }>\frac{ b }{ a }-\frac{ a }{ a }\] \[a*\ln \frac{ a }{ b }>b-a\] but this sint't going ight

OpenStudy (anonymous):

*this isn't going right